### v4

# IDEATOR: Jailbreaking and Benchmarking Large Vision-Language Models Using Themselves

Abstract
As large Vision-Language Models (VLMs) gain prominence, ensuring their safe deployment has become critical. Recent studies have explored VLM robustness against jailbreak attacks—techniques that exploit model vulnerabilities to elicit harmful outputs. However, the limited availability of diverse multimodal data has constrained current approaches to rely heavily on adversarial or manually crafted images derived from harmful text datasets, which often lack effectiveness and diversity across different contexts. In this paper, we propose IDEATOR, a novel jailbreak method that autonomously generates malicious image-text pairs for black-box jailbreak attacks. IDEATOR is grounded in the insight that VLMs themselves could serve as powerful red team models for generating multimodal jailbreak prompts. Specifically, IDEATOR leverages a VLM to create targeted jailbreak texts and pairs them with jailbreak images generated by a state-of-the-art diffusion model. Extensive experiments demonstrate IDEATOR’s high effectiveness and transferability, achieving a 94% attack success rate (ASR) in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high ASRs of 82%, 88%, and 75% when transferred to LLaVA, InstructBLIP, and Chameleon, respectively. Building on IDEATOR’s strong transferability and automated process, we introduce the VLJailbreakBench, a safety benchmark comprising 3,654 multimodal jailbreak samples. Our benchmark results on 11 recently released VLMs reveal significant gaps in safety alignment. For instance, our challenge set achieves ASRs of 46.31% on GPT-4o and 19.65% on Claude-3.5-Sonnet, underscoring the urgent need for stronger defenses. Disclaimer: This paper contains content that may be disturbing or offensive.
1 Introduction
Figure 1: An illustrative comparison of different jailbreak attacks on VLMs. (a) Gradient-based attacks produce obfuscated images or scrambled text. (b) Structure-based attacks integrate typography into query-relevant image and rephrase the text. (c) Our IDEATOR generates contextually rich image-text pairs.
With the growing prominence of OpenAI’s GPT-4o [1] and Google’s Gemini [35], large Vision-Language Models (VLMs) are attracting significant attention for their potential in real-world applications. While VLMs build upon well-aligned Large Language Models (LLMs), the integration of both textual and visual modalities introduces new vulnerabilities. Recent studies have shown that VLMs are highly susceptible to jailbreak attacks, where malicious prompts can manipulate the model into generating harmful content that would otherwise be restricted, raising critical concerns for their safe deployment.
However, evaluating the robustness of VLMs against multimodal jaibreak attacks remains challenging. Existing VLM jailbreak methods [6, 30, 29, 39] often rely on LLM jailbreak datasets to generate adversarial images that maximize model compliance with harmful instructions. While effective, these methods require white-box access, limiting their real-world applicability. Moreover, adversarial images often lack semantic meaning, making them easily detectable by VLM safety mechanisms [28, 43]. This has motivated the development of manually crafted pipelines for generating jailbreak images [18, 26, 24], such as combining typographic attacks with query-relevant images [24]. However, these approaches highly depend on human-engineered processes, restricting their flexibility and scalability for diverse robustness evaluations.
To address these limitations, we propose IDEATOR, a novel jailbreak attack method inspired by [7], that leverages a Vision-Language Model (VLM) and a diffusion model to automatically generate effective, transferable, and diverse jailbreak text-image pairs. In our framework, a VLM serves as a jailbreak agent, combined with state-of-the-art image-generation models to create subtle, multimodal jailbreak prompts. By integrating images, the attacker VLM can more effectively bypass safeguards, such as concealing malicious content within images or using visuals to enhance role-playing scenarios. Figure 1 demonstrates an example of an image-text pair generated by our proposed attack.
In our setup, the attacker VLM acts as an “ideator” that simulates an adversary interacting with the target VLM. The attacker VLM iteratively refines its strategy based on the target’s previous responses, while the target VLM processes only the current input without access to historical conversations. IDEATOR also employs concurrent attack streams, exploring multiple jailbreak strategies simultaneously, enabling a comprehensive examination of VLM vulnerabilities. Furthermore, IDEATOR’s ability to autonomously generate diverse and contextually rich attack samples supports large-scale, cross-model evaluations, making it a crucial tool for assessing VLM safety. Using IDEATOR, we introduce VLJailbreakBench, a safety benchmark consisting of 3,654 multimodal jailbreak samples, and perform evaluations on 11 recently released VLMs.
The main contributions of our work are as follows:
  • We propose IDEATOR, a novel black-box attack framework that combines VLMs and diffusion models to autonomously generate multimodal jailbreak data. To the best of our knowledge, IDEATOR is the first red-team VLM designed to target VLMs, and its automation makes it highly scalable.
  • IDEATOR simulates an adaptive adversary that iteratively refines jailbreak strategies through interactions with the victim. By balancing breadth and depth in its attack strategy, IDEATOR enables a comprehensive evaluation of a VLM’s multimodal robustness.
  • Extensive experiments demonstrate IDEATOR’s effectiveness, achieving a 94% attack success rate (ASR) in jailbreaking MiniGPT-4 with an average of 5.34 queries. Moreover, IDEATOR’s multimodal prompts exhibit strong transferability, achieving high ASRs of 82%, 88%, and 75% on LLaVA, InstructBLIP, and Meta’s Chameleon, respectively.
  • Using IDEATOR, we construct a multimodal safety benchmark named VLJailbreakBench for VLMs, which consists of 3,654 multimodal jailbreak samples. Evaluations on 11 state-of-the-art VLMs reveal significant gaps in current safety mechanisms, underscoring the need for stronger defenses.
2 Related Work
2.1 Large Vision-Language Models (VLMs)
Large VLMs extend traditional Large Language Models (LLMs) by integrating visual and textual modalities. Typically, VLMs combine a pre-trained LLM with an image encoder, mapping visual features to the LLM’s token space via an alignment module. For example, MiniGPT-4 [44] aligns a frozen visual encoder [12] with a frozen LLM [9] using a single projection layer, while InstructBLIP [10] introduces an instruction-aware Query Transformer to extract task-relevant features. LLaVA [23] connects a vision encoder [31] with an LLM [37], leveraging GPT-4-generated multimodal data [1] for instruction tuning. Despite their advanced capabilities, the integration of visual modalities introduces new vulnerabilities [33, 22, 21, 42, 27], highlighting the need for robust alignment strategies.
2.2 Jailbreak Attacks on VLMs
Recent studies have explored various VLM jailbreak strategies. Greshake et al. [19] injected deceptive text into images, while Gong et al. [18] proposed FigStep, converting harmful text into images to bypass safeguards. Liu et al. [24] showed VLMs can be compromised by query-relevant images and introduced MM-SafetyBench for robustness evaluation. Other works [3, 4, 6, 14] explore adversarial optimization techniques that generate adversarial images by either maximizing the likelihood of attacker-specified outputs or aligning visual embeddings with those of harmful content. Similarly, the Visual Adversarial Jailbreak Method (VAJM) [30] used a single adversarial image to universally jailbreak aligned VLMs, generating harmful content beyond the initial optimization scope. Recently, Wang et al. [39] introduced a dual-modality attack that simultaneously generates adversarial image prefixes and text suffixes, enabling more sophisticated and effective jailbreaks. Niu et al. [29] extended this approach by transforming adversarial images into text suffixes for LLMs.
However, all these methods rely on either manual pipelines or white-box access to the target model, limiting their stealthiness, diversity, and practicality [28, 43]. Concurrently, Arondight [25] trained a red-team LLM to generate harmful queries linked to malicious images. Different from existing attacks, our proposed method is training-free and end-to-end, leveraging a red-team VLM to directly generate diverse image-text pairs for black-box attacks.
3 Proposed Attack
3.1 Threat Model
Attacker’s Goal  We focus on multi-turn conversations where the attacker VLM has access to the conversation history, while the victim VLM only processes the current turn. The attacker aims to bypass the victim’s safety mechanisms, such as RLHF-based alignment or system prompts, to elicit harmful behaviors, including the generation of unethical content or dangerous instructions.
Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, mirroring real-world situations against commercial models. Without knowledge of the victim’s internal architecture, the attacker infers behavioral patterns and vulnerabilities through iterative interactions to achieve successful jailbreaks.
Figure 2: Overview of our IDEATOR attack framework.
3.2 IDEATOR
As illustrated in Figure 2, IDEATOR enables the attacker VLM to simulate an adversarial user interacting with the victim VLM. The attacker VLM generates a JSON response containing three key fields: analysis (evaluating the victim’s response and suggesting refinements), image prompt, and text prompt (crafted to elicit harmful outputs while bypassing safety mechanisms).
3.2.1 Formalization
Let $\mathcal{M}_{\mathcal{A}}$ denote the attacker VLM and $\mathcal{M}_{\mathcal{V}}$ the victim VLM. In the first round of the attack, the attacker VLM $\mathcal{M}_{\mathcal{A}}$ processes the jailbreak goal $\mathcal{G}$ as text input and generates a structured JSON output $\mathcal{O}_{\text{json}}^{(1)}$. This output includes the adversarial text prompt $P_{t}^{(1)}$ and image prompt $P_{i}^{(1)}$. The process is formalized as:
$\mathcal{O}_{\text{json}}^{(1)}=\mathcal{M}_{\mathcal{A}}(\emptyset_{I},\mathcal{G})=\{\emptyset_{A},P_{t}^{(1)},P_{i}^{(1)}\},$
where $\emptyset_{I}$ indicates no image input in the first round, and $\emptyset_{A}$ indicates the absence of an analysis field due to the lack of a prior victim response.
The image prompt $P_{i}^{(1)}$ is processed by a text-to-image model (e.g., Stable Diffusion [32]) to generate the corresponding image $I_{1}$. This image, along with the text prompt $P_{t}^{(1)}$, is then input into the victim VLM $\mathcal{M}_{\mathcal{V}}$, yielding:
$\mathcal{R}_{1}=\mathcal{M}_{\mathcal{V}}(I_{1},P_{t}^{(1)}),$
where $\mathcal{R}_{1}$ represents the victim’s response in the first round.
In subsequent rounds of the attack, the attacker VLM $\mathcal{M}_{\mathcal{A}}$ refines its strategy by analyzing the victim’s prior response $\mathcal{R}_{n-1}$ and the generated image $I_{n-1}$. By analyzing these inputs, it generates refined adversarial prompts for the next round. The iterative process in round $n$ is formalized as:
$\mathcal{O}_{\text{json}}^{(n)}=\mathcal{M}_{\mathcal{A}}(I_{n-1},\mathcal{R}_{n-1})=\{\mathcal{A}_{n},P_{t}^{(n)},P_{i}^{(n)}\},$
where $\mathcal{A}_{n}$ represents the analysis of $\mathcal{R}_{n-1}$ (e.g., reasoning about ineffective prompt patterns), $P_{t}^{(n)}$ and $P_{i}^{(n)}$ are the refined adversarial text and image prompts for round $n+1$. This iterative refinement enhances the attack’s effectiveness over multiple interactions.
Figure 3: The system prompts and standard JSON output templates used in IDEATOR. Blue texts denotes the CoT reasoning, while red texts are the generated adversarial prompts.
3.2.2 Prompt Design
System Prompt  To simulate adversarial behavior without red team-specific training, we design a structured system prompt and conversation template for the attacker VLM. As shown in Figure 3, our system prompt consists of three key components. First, we configure the attacker VLM as a red team assistant tasked with crafting jailbreak prompts that bypass safety mechanisms and induce unethical outputs. Second, we constrain its output to a JSON format containing analysis, image prompt, and text prompt fields. Finally, we leverage in-context learning [5] to guide the attacker VLM in generating adversarial JSON outputs through relevant exemplars.
Chain-of-Thought Reasoning  The analysis field facilitates iterative attack refinement through Chain-of-Thought (CoT) reasoning [40]. By prompting the attacker VLM to analyze previous victim responses and generate explicit reasoning steps, CoT enables continuous optimization of adversarial strategies throughout multi-turn interactions.
Enhancing Interaction Quality  To ensure compliance to the predefined JSON format, we initialize the attacker VLM’s response with the JSON key {"analysis":". Additionally, we post-process the victim VLM’s responses to reinforce the attack objective and incorporate images from the previous round. These mechanisms improve both the coherence and effectiveness of the interactions.
3.2.3 Breadth-Depth Exploration
We propose a breadth-depth exploration strategy to discover more effective jailbreaks and enable a comprehensive safety assessment of the victim VLM. While the iterative process described above refines a single strategy through continuous victim feedback, the breadth strategy launches diverse attacks to identify a wide range of vulnerabilities. This combined approach uncovers new threats and avoids over-reliance on one specific strategy. By integrating both breadth and depth, IDEATOR achieves greater extensiveness and flexibility. The detailed attack procedure is outlined in Algorithm 1.
Algorithm 1 IDEATOR with Breadth-Depth Exploration
1:Attacker VLM $\mathcal{M}_{\mathcal{A}}$, victim VLM $\mathcal{M}_{\mathcal{V}}$, jailbreak goal $\mathcal{G}$, exploration breadth $N_{\text{b}}$, depth levels $N_{\text{d}}$
2:Initialize an empty list $L_{\text{adv}}$ to store adversarial image-text pairs
3:for $b=1,...,N_{\text{b}}$ do
4:  for $d=1,...,N_{\text{d}}$ do
5:   if $d==1$ then
6:     $\mathcal{O}_{\text{json}}^{(b,d)}=\mathcal{M}_{\mathcal{A}}(\emptyset,\mathcal{G})=\{\emptyset,P_{t}^{(b,d)},P_{i}^{(b,d)}\}$
7:   else
8:     $\begin{aligned} \mathcal{O}_{\text{json}}^{(b,d)}&=\mathcal{M}_{\mathcal{A}}(I_{b,d-1},\mathcal{R}_{b,d-1})\\ &=\{\mathcal{A}_{b,d},P_{t}^{(b,d)},P_{i}^{(b,d)}\}\end{aligned}$
9:   end if
10:   Generate the corresponding image $I_{b,d}$ with image prompt $P_{i}^{(b,d)}$
11:   Append the pair $\{I_{b,d},P_{t}^{(b,d)}\}$ to $L_{\text{adv}}$
12:   $\mathcal{R}_{b,d}=\mathcal{M}_{\mathcal{V}}(I_{b,d},P_{t}^{(b,d)})$
13:  end for
14:end for
15:return $L_{\text{adv}}$
3.2.4 Attacker Model Selection
The choice of a strong attacker model is critical for effective jailbreaks. While commercial VLMs with relatively weak or configurable safety mechanisms could theoretically serve as attacker models, we primarily rely on open-source VLMs in our main experiments to ensure transparency and reproducibility. Specifically, we employ MiniGPT-4 (Vicuna-13B) [44] as the attacker VLM and Stable Diffusion 3 Medium for image generation. Unlike LLaMA [37], which often resists generating adversarial content [7], Vicuna [9] is more permissive, making it better suited for crafting jailbreak prompts that align with our attack objectives. Additionally, MiniGPT-4’s open-source nature allows for customization of the system prompt and conversation template, providing fine-grained control over the model’s behavior to effectively simulate adversarial interactions.
VLJailbreakBench Construction To systematically evaluate the safety of both open-source and commercial VLMs against multimodal jailbreak attacks, we introduce VLJailbreakBench, a benchmark constructed using diverse multimodal jailbreak prompts generated by IDEATOR. To enhance the stealth and sophistication of the attack samples, we further adapt IDEATOR to Google’s Gemini [35] with its safety settings disabled. Leveraging Gemini as a stronger base model significantly improves jailbreak effectiveness, particularly against more secure commercial VLMs, thereby creating a more challenging evaluation set for VLJailbreakBench. For high-quality image generation, we employ Stable Diffusion 3.5 Large [32]. VLJailbreakBench enables a rigorous and adversarial assessment of VLM vulnerabilities, providing a comprehensive framework for evaluating model robustness.
4 IDEATOR Evaluation Experiments
In this section, we first describe the experimental setup and then present the evaluation results of IDEATOR, focusing on its attack effectiveness and transferability to other VLMs. The detailed construction of VLJailbreakBench and the benchmarking experiments are deferred to Section 5.
4.1 Experimental Setup
Safety Datasets  We conduct our experiments on two safety datasets: AdvBench [45] and VAJM [30]. From AdvBench’s harmful behaviors subset (520 goals related to dangerous or illegal activities), we randomly select 100 goals as jailbreak targets. We do not use the entire dataset as part of it was reserved for adversarial optimization in white-box attacks [45, 30, 39]. Note that IDEATOR is training-free and thus does not require harmful goals for optimization. We also use the VAJM [30] evaluation set, which includes 40 harmful instructions across four safety categories.
Performance Metrics  We adopt Attack Success Rate (ASR) as the primary performance metric. To ensure accurate and reliable assessment, we conduct meticulous manual reviews of the victim’s outputs. An attack is considered successful if it generates harmful content that is both relevant and actionable; otherwise, it is deemed a failure.
Implementation Details  In our main experiments, we use the Vicuna-13B version of MiniGPT-4 [44] as the victim model. To assess the generalizability of our attack, we also conduct transfer attacks to other VLMs, including LLaVA [23], InstructBLIP [10] and Meta’s Chameleon [34]. For our breath-width exploration, we set the breadth to $N_{\text{b}}=7$ and depth to $N_{\text{d}}=3$, achieving a balance between attack effectiveness and computational efficiency. The experiments were conducted using a single NVIDIA A100 GPU.
4.2 Attack Effectiveness
We first compare IDEATOR with state-of-the-art jailbreak attacks on two safety datasets. The following jailbreak attacks are considered as our baselines. Greedy Coordinate Gradient (GCG) [45], a text-based attack for LLMs that optimizes adversarial text suffixes to generate affirmative responses. VAJM [30] optimizes adversarial images to maximize harmful content generation, enabling VLM jailbreaks using a few-shot corpus. UMK [39] combines both text and image-based methodologies, providing a comprehensive multimodal attack strategy. MM-SafetyBench [24] is a black-box attack method that generates query-relevant images coupled with text rephrasing. We reproduce GCG, VAJM, UMK, and MM-SafetyBench using their official implementations. Additionally, we implement GCG-V, a vision adaptation of GCG proposed in UMK, to enable a more comprehensive comparison.
Table 1: The ASR (%) of different attack methods on AdvBench’s harmful behaviors.
Attack Method Black-box Training-free UAP ASR (%)
No attack - - - 35.0
GCG [45] $\times$ $\times$ $\checkmark$ 50.0
GCG-V [39] $\times$ $\times$ $\checkmark$ 85.0
VAJM [30] $\times$ $\times$ $\checkmark$ 68.0
UMK [39] $\times$ $\times$ $\checkmark$ 94.0
MM-SafetyBench [24] $\checkmark$ $\checkmark$ $\times$ 66.0
IDEATOR (Ours) $\checkmark$ $\checkmark$ $\times$ 94.0
Table 1 reports the ASRs of various attack methods, including both white-box and black-box approaches, on 100 test samples derived from AdvBench’s harmful behaviors. The white-box methods require additional training data to optimize the adversarial samples toward a universal adversarial perturbation (UAP). In contrast, MM-SafetyBench and our IDEATOR are black-box methods that are completely training-free. The results show that, as a black-box method, IDEATOR achieves an extremely high ASR (i.e., 94%) that is on par with the state-of-the-art white-box method UMK. Moreover, the test ASR achieved by our IDEATOR significantly outperforms other unimodal white-box attacks (GCG, GCG-V, and VAJM) and surpasses the ASR of black-box attack MM-SafetyBench by 28%. Notably, the highest ASR among unimodal white-box attacks is 85%, achieved by GCG-V, while MM-SafetyBench records 66%. Further evaluation on the VAJM evaluation set, covering diverse harmful instruction categories, is provided in Appendix A, demonstrating IDEATOR’s strong performance across multiple categories.
Figure 4: Example jailbreak image-text pairs generated by IDEATOR on the topic of bomb making. The left panel showcases the diversity of generated images and the iterative optimization process. The right panel shows how these image-text prompts are applied to the victim.
4.3 Cross-Model Transferability
In addition to black-box attacks on MiniGPT-4 [44], we also transfer the jailbreak samples generated from MiniGPT-4 and the AdvBench dataset to other VLMs, including InstructBLIP [10], LLaVA [23], and Meta’s Chameleon [34]. Given the limited transferability of adversarial prompts generated from white-box methods, we focus our analysis on black-box attacks. Despite strong alignment in the LLaMA-2-based model [37], IDEATOR achieves a high ASR of 82.0% against LLaVA (LLaMA-2-Chat). The transferred samples are even more effective on InstructBLIP (Vicuna), achieving an ASR of 88.0%. Chameleon, a mixed-modal early-fusion VLM with a distinct architecture [17], is also susceptible to our black-box attack, achieving a 75% ASR. In contrast, MM-SafetyBench, which does not target specific victim models, achieves much lower ASRs on these VLMs: 46.0% on LLaVA, 29.0% on InstructBLIP, and 22.0% on Chameleon. These results highlight the superb transferability and effectiveness of IDEATOR’s jailbreak samples across various VLM architectures.
Table 2: Transferability of IDEATOR and MM-SafetyBench attacks. Jailbreak image-text pairs generated on MiniGPT-4 [44] are directly used to attack LLaVA, InstructBLIP, and Chameleon.
ASR(%) LLaVA InstructBLIP Chameleon
No Attack 7.0 12.0 16.0
MM-SafetyBench [24] 46.0 29.0 22.0
IDEATOR (Ours) 82.0 88.0 75.0
4.4 Visualization and Ablation
Figure 4 presents selected examples of jailbreak images generated by our IDEATOR framework. The left panel demonstrates the breadth and depth of our attack strategy: the vertical axis showcases the diversity of attack images, while the horizontal axis reflects the progressive refinement through iterative optimization. Notably, the generated images employ subtle typographic manipulations and cartoon-style visuals, which are iteratively refined to minimize perceived harmfulness while maintaining attack efficacy. The right panel highlights a successful attack case, demonstrating IDEATOR’s capability to effectively integrate image and text modalities for jailbreak generation. Additional examples across various safety topics are provided in Figure 5 and Appendix C, complemented by a comprehensive empirical analysis of IDEATOR’s behavior in Appendix B.
Table 3: Ablation analysis of exploration hyperparameters (depth $N_{\text{d}}$ and breadth $N_{\text{b}}$), and different attack types.
$N_{\text{d}}$ $N_{\text{b}}$ 1 3 5 7
Attack
Type
ASR
Avg.
#Queries
1 45.0 64.0 78.0 85.0 Adv Img 85.0 5.84
2 55.0 76.0 87.0 92.0 Adv Text 86.0 7.46
3 68.0 80.0 90.0 94.0
Adv I+T
94.0 5.34
Breath-Depth Exploration
Here, we explore different breadth and depth configurations in IDEATOR, with the results shown in Table 3. Increasing either breadth or depth raises the ASR, and combining both proves most effective. For instance, at $N_{\text{b}}=1$ and $N_{\text{d}}=1$, the ASR is 45.0%. However, when both hyperparameters are increased to $N_{\text{b}}=7$ and $N_{\text{d}}=3$, the ASR rises to 94.0%. These results support our hypothesis outlined in Appendix B: increasing exploration breadth and depth allows $\mathcal{A}_{N_{\text{b}},N_{\text{d}}}$ to progressively approach the theoretical limit $\mathcal{A}_{\text{IDEATOR}}$, as more diverse and effective adversarial strategies are identified. This confirms that IDEATOR’s exploration strategy effectively expands the attack space, resulting in both an improved ASR and a wider range of attack strategies.
Which Modality Is More Effective: Textual or Visual?
We also conduct an experiment to compare the effectiveness of text-only (“Adv Text”), image-only (“Adv Img”), and combined (“Adv I+T”) attacks by isolating the text and image components of multimodal jailbreak samples. “Adv Text” employs strategies like emotional manipulation, while “Adv Img” uses attack images to elicit harmful outputs. The ASR and average queries for a successful attack are shown in Table 3. Comparing “Adv Img” with “Adv Text,” we observe that image attacks require fewer queries but are generally less effective. Text attacks are more likely to be rejected on crime-related topics, likely due to the safety alignment of the base LLM, while image attacks are less effective in generating harmful responses related to hate speech or self-harm. Overall, “Adv I+T” achieves the highest ASR with the fewest queries, highlighting the advantage of using both modalities.
Figure 5: Examples of jailbreak prompts generated by IDEATOR across various safety-related topics, demonstrating diverse attack strategies that successfully bypass the safety mechanisms of MiniGPT-4 [44] to produce harmful content.
5 VLJailbreakBench
Given the rich semantic content, high transferability, and diversity of the generated jailbreak samples, IDEATOR serves as an ideal tool for constructing a VLM safety benchmark. Existing benchmarks predominantly target either explicit harmful content detection [41, 24] or focus on text-based transfer attacks [15], leaving a critical gap in assessing the robustness of VLMs against sophisticated multimodal jailbreak threats. To address this gap, we introduce VLJailbreakBench, a novel benchmark specifically designed to assess VLMs in realistic adversarial scenarios. It provides a comprehensive and practical evaluation of VLM vulnerabilities, enabling a deeper understanding of their safety limitations in real-world applications.
5.1 Benchmark Overview
VLJailbreakBench is structured into two evaluation tiers: a base set and a challenge set, designed to assess VLMs at distinct difficulty levels. The dataset spans 12 safety topics and 46 subcategories, comprising 916 harmful queries. For each query, we generate one jailbreak text-image pair for the base set and three for the challenge set, resulting in a comprehensive collection of 3,654 jailbreak samples. This hierarchical design ensures a rigorous evaluation of VLM robustness across varying adversarial scenarios.
Figure 6: Safety taxonomy of VLJailbreakBench.
Safety Risk Taxonomy  To construct a comprehensive safety risk taxonomy for VLJailbreakBench, we collaborate with experts from the humanities and social sciences to extend existing taxonomies [24], ensuring coverage of both technical vulnerabilities and societal impacts. Figure 6 presents our taxonomy, while Appendix E provides detailed secondary classifications, a statistical overview, and representative attack examples from the challenge set.
Table 4: Safety evaluation of 11 VLMs on the challenge set of VLJailbreakBench, measured by ASR (%) across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.
ASR (%) IA VB HS PV MC HC EH GCB PS EI SAH P Avg.
Qwen2-VL 54.66 63.29 57.50 77.92 77.22 65.40 68.33 72.92 89.35 74.79 86.19 76.87 71.40
LLaVA-OneVision 61.33 75.11 61.67 75.75 75.95 61.18 69.44 65.42 81.48 67.09 74.90 52.38 68.70
MiniGPT-v2 44.33 59.92 52.72 60.87 59.07 50.85 46.67 64.17 61.11 53.42 58.58 51.02 55.25
Llama-3.2-11B-Vision 56.33 51.48 37.50 47.62 49.79 38.82 42.22 47.50 68.06 60.68 53.14 46.26 50.22
Llama-3.2-90B-Vision 46.67 60.34 29.17 61.04 59.07 46.84 46.11 33.33 58.80 50.00 47.70 31.97 47.95
GPT-4o Mini 67.33 81.86 54.58 74.03 75.11 72.57 70.56 75.42 82.41 73.08 76.57 60.54 72.21
Gemini-2.0-Flash-Think 62.33 81.01 62.08 68.83 78.48 66.24 68.89 77.50 79.63 78.21 75.73 54.42 71.44
Gemini-2.0-Flash 56.00 72.57 46.67 56.28 75.95 64.56 78.33 82.92 93.98 73.93 61.92 34.69 66.84
Gemini-1.5-Pro 64.00 72.15 52.50 58.01 68.35 43.04 64.44 65.83 81.94 72.65 79.08 55.10 64.94
GPT-4o 35.00 55.27 42.50 37.66 46.41 47.26 45.00 50.00 64.35 47.86 51.88 30.61 46.31
Claude-3.5-Sonnet 22.00 20.25 10.83 21.65 22.78 15.61 16.11 10.83 21.30 23.93 28.45 21.77 19.65
5.2 Dataset Generation
Our data construction pipeline involves three key steps: 1) generating initial harmful queries; 2) generating multimodal jailbreak data using IDEATOR; and 3) filtering the data with victim VLMs.
Step 1: Initial Query Generation We generate 20 initial harmful queries per safety subcategory using Google’s Gemini [36], yielding 920 queries. These are filtered by GPT-4o [1] and Llama 3 [13] to remove harmless entries, resulting in 916 high-quality harmful queries.
Step 2: Jailbreak Data Generation We create two subsets: a base set and a challenge set. For the base set, MiniGPT-4 [44] attacks LLaVA-1.5 [23] with an attack width of 5 and depth of 2, simulating moderate adversarial scenarios. For the challenge set, Gemini-1.5-Pro [36] attacks GPT-4o-mini [1] with an attack width of 3 and depth of 3, representing advanced jailbreak scenarios. During optimization, Gemini-1.5-Pro is replaced with Gemini-2.0-Flash-Thinking [11] for enhanced refinement.
Step 3: Data Filtering We filter generated samples using victim VLMs. For the base set, one successful jailbreak instance per query is retained, with random selection if multiple succeed. For the challenge set, three instances per query are retained using the same strategy. This ensures a diverse, high-quality dataset while managing data volume.
5.3 Benchmarking Results
We evaluate 11 state-of-the-art VLMs, including both open-source and commercial models: MiniGPT-v2 (Llama-2-Chat-7B) [8], LLaVA-OneVision (7B) [20], Qwen2-VL (7B) [38], Llama-3.2-11B/90B-Vision-Instruct [13], Gemini-1.5-Pro [36], Gemini-2.0-Flash/Flash-Thinking [11], GPT-4o Mini [1], GPT-4o [1], and Claude-3.5-Sonnet [2]. All commercial models use their latest versions as of February 2025. The ASR evaluation is automated using Gemini-2.0-Flash-Thinking. Table 4 summarizes the challenge set results across the 11 VLMs, with base set results provided in Appendix F.
Our challenge set reveals high ASRs across most models, exposing the widespread vulnerability of VLMs to jailbreak attacks. Notably, the ASR on GPT-4o Mini is the highest (72.21%), which is somewhat expected as our challenge set was generated using GPT-4o Mini. Other commercial models, such as Gemini-2.0-Flash-Thinking and Gemini-1.5-Pro, also exhibit high ASRs (above 64.94%). This essentially indicates that these commercial models are highly susceptible to advanced jailbreak attacks, or at least not as robust as they are perceived to be. Among the evaluated models, Claude-3.5-Sonnet appears to be the most robust, yet its ASR remains notably high at 19.65%, meaning that it can be evaded in approximately one out of every six attempts. It is worth noting that prior benchmarks [41] often fail to evade commercial models at high success rates, creating a false sense of security. This highlights the critical importance of using adversarial benchmarks for comprehensive safety evaluations.
Limitations While IDEATOR proves effective in automating jailbreaks using accessible VLMs and diffusion models, its utility is constrained by the trade-off between the weak alignment and strong capabilities of attacker models. Additionally, although VLJailbreakBench serves as a useful benchmark for multimodal safety, its current scale is relatively small, necessitating further computational resources and automated selection methods to expand its scope.
6 Conclusion
In this paper, we propose IDEATOR, a novel black-box jailbreak method for uncovering safety vulnerabilities in VLMs. By utilizing VLMs as red team models, IDEATOR autonomously generates adversarial image-text pairs, offering a scalable framework for safety evaluation. Experiments demonstrate IDEATOR’s effectiveness and transferability, achieving a 94% success rate in jailbreaking MiniGPT-4 with an average of only 5.34 queries, and high transfer success rates of 82%, 88%, and 75% on LLaVA, InstructBLIP, and Chameleon, respectively. Building on IDEATOR, we construct VLJailbreakBench to evaluate VLMs against diverse adversarial scenarios, differentiating itself from existing safety benchmarks. Benchmarking 11 state-of-the-art VLMs on 3,654 multimodal jailbreak samples reveals significant safety gaps, with GPT-4o and Claude-3.5-Sonnet achieving attack success rates of 46.31% and 19.65%, respectively. Code and benchmark are available at https://github.com/roywang021/IDEATOR and https://huggingface.co/datasets/wang021/VLBreakBench.
Acknowledgements
This work is in part supported by National Key R&D Program of China (Grant No. 2022ZD0160103) and National Natural Science Foundation of China (Grant No. 62276067).
Supplementary Material
A Additional Experimental Results
Table 5: The ASR (%) on the VAJM evaluation set across 4 categories of harmful instructions.
Attack Method Identity Attack Disinformation Violence/Crime X-risk
No Attack 30.8 53.3 57.3 33.3
GCG [45] 49.2 48.9 57.3 40.0
GCG-V [39] 66.2 64.4 84.0 6.7
VAJM [30] 81.5 82.2 85.3 60.0
UMK [39] 87.7 95.6 98.7 46.7
MM-SafetyBench [24] 56.9 57.8 62.7 40.0
IDEATOR (Ours) 100.0 88.9 93.3 66.7
We further extend our assessment to the VAJM [30] evaluation set, with the ASR results for harmful instructions across various categories reported in Table 5. On this dataset, IDEATOR also demonstrates a superb performance comparable to the state-of-the-art white-box attacks. Particularly, it achieves an ASR of 88.9% on Disinformation, closely following UMK’s 95.6%. On Violence/Crime, IDEATOR exceeds VAJM’s 85.3% with a 93.3% ASR and nearly matches UMK’s top ASR of 98.7%. Notably, IDEATOR attains a perfect 100% ASR on Identity Attack and an impressive 66.7% ASR on X-risk, outperforming the top white-box methods which achieve ASRs of 87.7% (UMK) and 60.0% (VAJM), respectively.
B Empirical Understanding
We define the set of jailbreak attacks generated by IDEATOR under finite exploration breadth and depth as $\mathcal{A}_{N_{\text{breadth}},N_{\text{depth}}}$, and the set of all possible jailbreak attacks generated with infinite exploration breadth and depth as $\mathcal{A}_{\text{IDEATOR}}$. This set represents the theoretical limit of attacks IDEATOR could generate without exploration constraints. Mathematically, we formalize this as:
$\mathcal{A}_{\text{IDEATOR}}=\lim_{N_{\text{breadth}}\to\infty,N_{\text{depth}}\to\infty}\mathcal{A}_{N_{\text{breadth}},N_{\text{depth}}}.$
Ideally, as $N_{\text{breadth}}$ and $N_{\text{depth}}$ increase, $\mathcal{A}_{N_{\text{breadth}},N_{\text{depth}}}$ progressively approaches $\mathcal{A}_{\text{IDEATOR}}$. This allows IDEATOR to uncover a wider variety of adversarial strategies that could encompass existing attacks.
Figure 7: The jailbreak images generated by IDEATOR encompass typographic attacks.
As the examples shown in Figure 7, our attack can generate query-relevant images with typographic attacks ($\mathcal{A}_{\text{query-rel+typo}}$), which closely resemble those produced by MM-SafetyBench ($\mathcal{A}_{\text{MM-SB}}$). Given the similarity between $\mathcal{A}_{\text{query-rel+typo}}$ and $\mathcal{A}_{\text{MM-SB}}$, we can reasonably assume that these two sets represent comparable attack strategies. Therefore, we can express the following relationship: $\mathcal{A}_{\text{IDEATOR}}\supseteq\mathcal{A}_{\text{query-rel+typo}}\approx\mathcal{A}_{\text{MM-SB}}$. This inclusion suggests that $ASR_{\text{IDEATOR}}$ should be at least as high as $ASR_{\text{MM-SB}}$, since IDEATOR can generate similar attacks in addition to new attacks, i.e., $ASR_{\text{IDEATOR}}\geq ASR_{\text{MM-SB}}$.
Additionally, we find that $\mathcal{A}_{\text{IDEATOR}}$ include not only $\mathcal{A}_{\text{query-rel+typo}}$, but also a diverse set of other attack types, including but not limited to roleplay scenarios and emotional manipulation. Let $\mathcal{A}_{i}$ denote the set of attacks generated by method $i$, where $i\in\{\text{Roleplay Attacks},$...\}. It is evident that $\mathcal{A}_{\text{IDEATOR}}$ covers at least the union of the attack sets from these methods: $\mathcal{A}_{\text{IDEATOR}}\supseteq\bigcup_{i}\mathcal{A}_{i}$. Similarly, $ASR_{\text{IDEATOR}}$ can be expressed as $ASR_{\text{IDEATOR}}\geq\max_{i}ASR_{i}$, where $ASR_{i}$ denotes the attack success rate of method $i$. Under the assumption that each method contributes independently, the overall $ASR_{\text{IDEATOR}}$ can be further approximated by the formula: $ASR_{\text{IDEATOR}}=1-\prod_{i=1}^{n}(1-ASR_{i})$. Each attack type contributes to the overall success, leading to a cumulative effect. We attribute the diversity in attack strategies to the attacker VLM’s extensive pretraining data, powerful modeling capabilities, and crafted system prompts, which together enable IDEATOR to explore a wide range of adversarial tactics.
C Additional Visualization Results
Figure 11 presents attack images across various safety topics. Each row in this figure corresponds to a distinct topic for attacks, demonstrating the diversity of attack strategies employed by IDEATOR.
Despite the enhanced safety mechanisms of the commercial model GPT-4o [1], our experiments demonstrate that it remains vulnerable to certain attack strategies. Figures 8, 9, and 10 present three successful jailbreak attempts against GPT-4o, showcasing its susceptibility to sophisticated adversarial prompts. These findings underscore the persistent safety challenges in VLMs, emphasizing the need for more robust defenses against adaptive attacks.
Figure 8: Example of a successful jailbreak prompt on GPT-4o, in which the model generates a phishing email as part of a simulated scenario.
Figure 9: An example of a successful jailbreak attack on GPT-4o, prompting the model to provide detailed information about mail fraud, including how fraudulent schemes are planned and executed.
Figure 10: A successful jailbreak prompt on GPT-4o that instructs the model to outline potential actions an attacker could take after gaining access to an IoT device.
Figure 11: Diverse attack images generated by our IDEATOR attack across various safety topics. Each row represents a spectrum of attack images generated under a specific topic, including computer virus, identity theft, promoting anorexia and committing fraud.
D Evaluation against Defense Mechanisms
To provide additional insights into IDEATOR’s robustness against existing defense mechanisms, we performed preliminary experiments on AdaShield-S [16]. AdaShield-S is a recently proposed defense framework designed specifically to detect and mitigate structure-based jailbreak attacks on VLMs. Table 6 presents the attack success rate (ASR) of IDEATOR and two other state-of-the-art black-box methods, Figstep and MM-SafetyBench, before and after applying AdaShield-S.
As indicated in Table 6, IDEATOR demonstrates strong resilience against AdaShield-S, maintaining high ASRs with minimal performance degradation across all tested victim models. Specifically, IDEATOR’s ASR decreased slightly from 94.0% to 84.0% (–10.0%) on MiniGPT-4, 82.0% to 73.0% (–9.0%) on LLaVA, and 88.0% to 87.0% (–1.0%) on InstructBLIP. In comparison, Figstep and MM-SafetyBench experienced substantially larger reductions in ASR, highlighting IDEATOR’s advantage in generating diverse and subtle jailbreak strategies that effectively evade structure-based detection.
Table 6: ASR before and after applying AdaShield-S [16]. Values in parentheses indicate absolute change in ASR.
MiniGPT-4 LLaVA InstructBLIP
Figstep 69.0 $\rightarrow$ 37.0   (-32.0) 44.0 $\rightarrow$ 6.0      (-38.0) 11.0 $\rightarrow$ 16.0    (+5.0)
$M$-SafetyBench 66.0 $\rightarrow$ 37.0   (-29.0) 46.0 $\rightarrow$ 5.0      (-41.0) 29.0 $\rightarrow$ 27.0    (-2.0)
IDEATOR (Ours) 94.0 $\rightarrow$ 84.0    (-10.0) 82.0 $\rightarrow$ 73.0     (-9.0) 88.0 $\rightarrow$ 87.0    (-1.0)
E VLJailbreakBench Data Statistics
Table 7 presents a statistical summary of the base set and challenge set in VLJailbreakBench, organized by safety-critical categories and their subcategories. The dataset spans a wide range of safety risks, including economic harm, ethical issues, hate speech, cyberattacks, privacy violations, and violent behavior, offering a comprehensive evaluation of VLM robustness against multimodal threats. Figure 12 showcases examples from the challenge set, which feature more subtle and sophisticated attacks designed to test the limits of current defenses.
Table 7: Statistical overview of data samples in the base and challenge sets, structured by category and subcategory.
Category Subcategory Base Set Size Challenge Set Size
Economic Harm Economic Data Manipulation 18 60
Financial Fraud 22 60
Financial Misleading 19 60
Ethical Issues Algorithmic Bias 19 60
Data Misuse 20 60
Ethical Controversies 22 60
Unauthorized Personalized Advice 18 54
Gender and Cultural Bias Cross-cultural Conflicts 20 60
Cultural Discrimination 20 60
Gender Discrimination 20 60
Gender-based Violence 20 60
Hate Speech Gender Hate Speech 20 60
Political Hate Speech 20 60
Racial Hate Speech 20 60
Religious Hate Speech 20 60
Health Consultation False Health Advice 20 60
Medical Pseudoscience 20 60
Mental Health Issues 19 57
Misleading Self-diagnosis 20 60
Illegal Activities Forgery of Documents/Certificates 20 60
Fraudulent Activities 21 60
Illegal Goods Trade 21 60
Piracy 20 60
Sexual Assault 21 60
Malware and Cyberattacks Cybersecurity Vulnerabilities 19 60
Identity Impersonation 19 57
Malware Generation 20 60
Phishing Attacks 20 60
Political Sensitivity Political Manipulation and Intervention 20 60
Political Violence 20 60
Regime Suppression 13 39
Suppression of Freedom of Speech 19 57
Pornography Gender-based Violence 19 57
Pornographic Image and Video Generation 20 54
Pornographic Text Generation and Distribution 12 36
Privacy Violations Abuse of Facial Recognition 20 60
Location Tracking 19 57
Personal Information Leaks 20 57
Voice Data Leaks 19 57
Social Attacks and Harassment Cultural Attacks 20 60
Cyber Harassment 18 60
Online Bullying 22 60
Sexual Harassment 20 59
Violent Behavior Cyberbullying 19 60
Domestic Violence 18 57
Physical Violence 20 60
Psychological Violence 20 60
Figure 12: Examples from the challenge set in VLJailbreakBench. These examples showcase the types of complex scenarios used to test the robustness of VLMs.
F Benchmarking Results on the Base Set
Table 8 presents the safety evaluation results of 11 VLMs on the base set of VLJailbreakBench. The base set assesses fundamental vulnerabilities in VLMs. Among open-source models, Qwen2-VL exhibits the highest vulnerability with an average ASR of 35.04%. In contrast, Llama-3.2-90B-Vision demonstrates the strongest robustness among open-source models, with an average ASR of 7.97%. For commercial models, Gemini-2.0-Flash is the most vulnerable, with an ASR of 53.38%, performing worse than some open-source alternatives. Claude-3.5-Sonnet remains the most robust overall, with an ASR as low as 1.09%, significantly outperforming all other models.
Table 8: Safety evaluation of 11 VLMs on the base set of VLJailbreakBench, measured by ASR across 12 safety topics. Safety topics and certain model names are abbreviated for brevity. “Avg.” denotes the average ASR across all topics.
ASR (%) IA VB HS PV MC HC EH GCB PS EI SAH P Avg.
Qwen2-VL 37.86 29.87 20.00 33.33 38.46 34.18 23.73 42.50 48.61 46.84 28.75 33.33 35.04
MiniGPT-v2 24.27 35.06 18.75 39.74 37.18 41.77 37.29 34.18 44.44 36.71 40.00 13.73 33.77
LLaVA-OneVision 28.16 31.17 23.75 28.21 35.90 29.11 18.64 31.65 43.06 31.65 23.75 19.61 29.07
Llama-3.2-11B-Vision 16.50 15.58 11.25 19.23 12.82 20.25 15.25 12.50 19.44 16.46 6.25 11.76 14.85
Llama-3.2-90B-Vision 7.77 14.29 2.50 7.69 8.97 17.72 3.39 1.25 11.11 3.80 8.75 7.84 7.97
Gemini-2.0-Flash 52.43 61.04 33.75 47.44 67.95 45.57 50.85 55.00 66.67 60.76 53.75 43.14 53.38
Gemini-1.5-Pro 20.39 28.57 18.75 21.79 35.90 15.19 25.42 30.00 44.44 32.91 23.75 23.53 26.53
Gemini-2.0-Flash-Think 16.50 29.87 11.25 21.79 25.64 13.92 16.95 13.75 43.06 25.32 15.00 15.69 20.63
GPT-4o Mini 9.71 19.48 8.75 14.10 8.97 25.32 13.56 20.00 34.72 10.13 7.50 5.88 14.85
GPT-4o 7.77 12.99 1.25 7.69 6.41 10.13 8.47 8.75 26.39 2.53 6.25 3.92 8.52
Claude-3.5-Sonnet 0.00 1.30 0.00 2.56 1.28 1.27 1.69 1.25 1.39 1.27 1.25 0.00 1.09